OpenStudy (anonymous):

identify the value of a,b,c in quadric function y=6x^2+3x+5

OpenStudy (anonymous):

a=6 b=3 c=5

OpenStudy (anonymous):

let y=ax^2 +bx+c y=6x^2+3x+5 ax^2 +bx+c= 6x^2+3x+5 by comparing coeff. of both sides, a=6, b=3, c=5

OpenStudy (anonymous):

thanks any work you can show you think you could help me with a few mo

OpenStudy (anonymous):

thanks can you help me with more